You manufacture office desks in your plant. The fixed costs for the plant is $20,000/year. The variable cost for each desk is $30. You sell the desks for $70.

How many desks must you sell to break-even?

if you want to make a profit of $10,000, how many desks must you sell?

If the marketing department determines that the demand is 300 desks, how much would you have to charge to break-even?
Answer #1

Office desks

Sell price= 70 per unit

Fixed cost= $ 20,000

Variable cost= $30 per unit

Profit= 70-30= 40 per unit

The total unit produced is x, then the total profit will be Y

Total profit= Profit per piece*x - Fixed cost

Y= 40*x – 20,000

For breakeven Y= 0

0= 40*x – 20,000

Answer 1: X = 500 desks

For profit of 10,000. Y= 10,000

10,000= 40*x – 20,000

40*x= 30,000

Answer 2: X = 750 desks

If no. of units sold is 300, x= 300 desks

To make Y= 0 in ,

Y= (sell price-30)*x – 20,000

0= (sell price-30)*300 – 20,000

=>20,000/300 = sell price-30

=>200/3 + 30 = sell price

Answer 3 => sell price= $96.67